A boat took 1 hour and 50 minutes to go 55 miles downstream and 3 hours and 40 minutes to return. What is the rate of the current?
Answer by mananth(16949)

against current 3.67 hours
with current 1.83 hours

Distance = same=55
boat speed =x
current speed =y
t=d/r
55/(x-y)=3.67
3.67(x-y)=55.00
3.67x-3.67y= 55 ....................1
55/(x+y)=1.83
1.83(x+y)=55
1.83x+1.83y= 55 ...............2
Multiply (1) by 1.83
Multiply (2) by 3.67
we get
6.72x-6.72y= 100.83
6.73x+6.73y= 201.85
13.45 x=302.68
/13.45
x=22.5mph boat speed

plug value of x in (1)
3.67x-3.67y= 55
82.51-3.67y= 55
-3.67 y= 55 -82.51
-3.67 y= -27.51
y= 7.5 mph current speed